OpenStudy (anonymous):

I'm trying to do the third plot of 3x-9y=3. Can anyone help me?

OpenStudy (anonymous):

3y=x-1

OpenStudy (anonymous):

Thanks!!!

OpenStudy (anonymous):

I'm sorry but I didn't even finish answering the question

OpenStudy (anonymous):

what do you mean by third plot

OpenStudy (anonymous):

I'm trying to graph 3x-9y=9

OpenStudy (anonymous):

take x=0 then y=-1 then take y=0 then x=3 so the coordinates are (0,-1) and (3,0)

OpenStudy (anonymous):

so mark those two points. draw a line going through those two points. and you have your graph

OpenStudy (anonymous):

Wow thanks......

OpenStudy (masumanwar):

simplify inters of y as a function of x as y=x/3-1/3 which represents a straight line and then plot the graph in range of values defined otherwise any arbitrary value within a suitable range